OpenStudy (anonymous):

A hotel charges an 8% tax, as required by law to pay the city. If Mr. and Mrs. Lafferty paid $167.40, what was their bill before the tax was added to it?

OpenStudy (anonymous):

x = 155 solve equation: 1.08 * x = 167.4 x = 167.4/ 1.08 x = 155

OpenStudy (anonymous):

See similar calculation on percentage calculator: If a goods sells for 167.4 after a -8% markdown, what was its original price? ................ 155 http://www.hackmath.net/en/calculator/percentage?a1=&b1=&a2=&b2=&a3=&b3=&submit1=Calculate&a4=&b4=&a5=&b5=&a6=&b6=&a7=&a8=&b8=&b9=&a10=167.4&b10=-8